A six sided die has been rolled 3 times to randomly define three digits of a code. We know that the code is not 111 or 666. What is the chance of finding the code in the first attempt?
Answer #1

Solution:

A die has total six outputs from 1 to 6.

So, number of three digit codes = 6*6*6 = 216

We are given that the code is not 111 or 666.

So, number of available codes = 216 – 2 = 214

Probability of finding the code in the first attempt = 1/214 = 0.004673

Required probability = 0.004673